OpenStudy (anonymous):

What is the largest odd number that is a factor of 860 ?

OpenStudy (anonymous):

\[860=2^2\times 5\times 43\]

OpenStudy (anonymous):

no the answer is 215

OpenStudy (anonymous):

that's what i though too but it isn't

OpenStudy (hunus):

I think he was giving you a hint of where to go. You can get the factors from the prime factorization The factors are: 2 4 5 10 20 86 172 215 You get these by multiplying the numbers in your prime factorization

OpenStudy (anonymous):

oh thx

OpenStudy (hunus):

Yup :)